OpenStudy (anonymous):

use substitution to solve system of equations. y=x+5 3x+y=25

OpenStudy (anonymous):

x=5, y=10

OpenStudy (anonymous):

1. 3X+Y=25 2. Y= X+5 Since Y is already solved for it is in Slope Intercept form ( Y=mx+b) plug that value in to # 1's Y value: 3X+Y=25 3X +( X+5)= 25 3X + X+ 5=25 Combine like terms: 4X + 5 = 25 To solve the equation get variables on one side and numbers on the other by subtracting -5 from each side: 4X +5-5=25-5 4x=20 Divide by each side to get 4: X= 5 Now Back Substitute to get the other point in the answer, plug in X=5 into Y= X+5 and solve: Y= X+ 5 Y= (5)+5 Y= 10 Solution is: (5, 10)

OpenStudy (anonymous):

y=x+5 , 3x-25=-y, make an addition you will get ,4x=20 so x=20/4 which is 5, y=x+5 so then ,replace by the value of x , y=10

OpenStudy (anonymous):

Y=4x+5 2x+y=17 (2,13) how do i use substitution to solve system of equations?